Esther, Vida, and Clair were asked to consider two different cash flows: GH¢1000 that they could receive today and GH¢3000 that would be received 3 years from today. Esther wanted the GH¢1000 today, Vida chose to collect GH¢3000 in 3 years, and Clair was indifferent between these two options.
1
Expert's answer
2020-06-29T19:04:10-0400

"\text{Esther made the right option which gives her maximum value at the end of 3 years than Vida and Clair.}""\text{ Esther has made a decision to collect GH¢1000 now. }"

"\text{To calculate future value of Esther choice at the end of 3 year we will use formula below}"

"FV = PV (1 + r)^{t}"

"\text{Discount at year 3 is 10}"% "\text{ is 0.751}"

"\text{ Future value will be ;}"

"FV = GH\u00a22487."

"\text{Vidha choose to collect GH¢3000 at end of 3 year}"

"\text{The present value of GH¢3000 at 10}" %"\text{ discount rate will be;}"

"PV = FV (1 + r)^{-t}"

"PV = 3000 (0.751)"

"PV = GH\u00a22253."

"\text{Clair is indifferent from the choices given but this is also not a smart choice.}"

"\text{Esther has made the right choice among the three women , }""\text{she gets maximum value}" of"\text{ GH¢2487.}"